OpenStudy (ell):

Write the exponential expression with only positive exponents. 3 * x^-2

OpenStudy (anonymous):

hint : turn it into a fraction.

OpenStudy (jim766):

to change a variable with a neg exponent to positive exponent take its reciprocal (flip the fraction ) and make the exponent positive \[3* x^-2 = 3*\frac{ 1 }{ x^2 }= \frac{ 3 }{ x^2 }\] make sense?

OpenStudy (ell):

it does, thanx! thanx @timo86m

OpenStudy (ell):

So would you multiply straight across?

OpenStudy (jim766):

yes, 3 is really 3/1, when multiplying fractions just go straight across

OpenStudy (anonymous):

use this rule